# Asymmetries and Circumstellar Interaction in the Type II SN 2024bch

### v1

@@ 7 Summary and Conclusions
We have presented comprehensive photometric and spectroscopic data of the Type II SN 2024bch from the UV to IR over the first 150 days. The 2.9 day optical spectra shows lines of H I He II, C IV, and N IV but lack C III, N III, and O IV. By day 6 all but the narrow Balmer lines persist. High-resolution spectra reveal narrow P-Cygni in H that constrains the progenitor wind to 35–40 km s-1, speeds typically associated with RSGs. Comparison between models and the early spectra indicate a mass loss rate of between . This requires a period of enhanced mass in the years to weeks before explosion as these values are four orders of magnitude higher than normal quiescent winds from RSGs.
After day 60 when the lightcurve begins to fall from the plateau we also see the emergence of multi-peaked H, likely from the SN shock still interacting with an asymmetric CSM (Smith et al., 2015). The lack of a corresponding red component, and the persistence of narrow H absorption points to CSM in a disc or torus that we are viewing nearly edge-on. This CSM is an order of magnitude further out than the CSM giving rise to early interaction signatures, which hints at a complex CSM geometry and mass-loss history for the progenitor of SN 2024bch. Additionally, the H I and [O I] lines all show blueshifted peaks that persist well into the nebular phase, likely from asymmetries in the SN explosion itself.
@@ Acknowledgements
J.E.A. is supported by the international Gemini Observatory, a program of NSF’s NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. Based on observations obtained at the international Gemini Observatory (GN-2023B-Q-135, GN-2024A-Q-403), a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation on behalf of the Gemini Observatory partnership. Time-domain research by the University of Arizona team and D.J.S. is supported by National Science Foundation (NSF) grants 2108032, 2308181, 2407566, and 2432036 and the Heising-Simons Foundation under grant #2020-1864. S.V. and the UC Davis time-domain research team acknowledge support by NSF grants AST-2008108.
@@
| 2024-01-31 | 60340.45 | 2.9 | Gemini-N+GMOS | 5800 | 600 4 |
| 2024-01-31 | 60340.46 | 2.9 | FTN+FLOYDS | 380 | 1800 |
| 2024-02-01 | 60341.48 | 3.9 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-02-03 | 60343.54 | 6.0 | FTN+FLOYDS | 380 | 900 |
| 2024-02-04 | 60344.48 | 6.9 | FTN+FLOYDS | 380 | 900 |
| 2024-02-04 | 60344.62 | 7.1 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-02-05 | 60345.46 | 7.9 | FTN+FLOYDS | 380 | 900 |
| 2024-02-05 | 60345.52 | 8.0 | LBT+PEPSI | 50000 | 5400 |
| 2024-02-06 | 60346.32 | 8.8 | FTN+FLOYDS | 380 | 600 |
| 2024-02-09 | 60349.57 | 12.0 | FTN+FLOYDS | 380 | 600 |
| 2024-02-11 | 60351.30 | 13.8 | FTN+FLOYDS | 380 | 600 |
| 2024-02-13 | 60353.30 | 15.8 | Bok+B&C | 700 | 600 3 |
| 2024-02-14 | 60354.37 | 16.8 | FTN+FLOYDS | 380 | 600 |
| 2024-02-16 | 60356.65 | 19.1 | Xinglong+BFOSC | 400 | 2100 |
| 2024-02-19 | 60359.34 | 21.8 | FTN+FLOYDS | 380 | 600 |
| 2024-02-19 | 60359.43 | 21.9 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-02-22 | 60362.40 | 24.9 | FTN+FLOYDS | 380 | 600 |
| 2024-02-25 | 60365.49 | 28.0 | FTN+FLOYDS | 380 | 600 |
| 2024-02-28 | 60368.35 | 30.8 | FTN+FLOYDS | 380 | 600 |
| 2024-02-29 | 60369.41 | 31.9 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-03-02 | 60371.10 | 33.6 | Bok+B&C | 700 | 600 3 |
| 2024-03-02 | 60371.27 | 33.7 | FTN+FLOYDS | 380 | 600 |
| 2024-03-11 | 60380.23 | 42.7 | LBT+PEPSI | 50000 | 14400 |
| 2024-03-17 | 60386.51 | 49.0 | FTN+FLOYDS | 380 | 600 |
| 2024-03-20 | 60389.30 | 51.8 | Lick+KAST | 800 | 1200 |
| 2024-03-24 | 60393.43 | 55.9 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-03-25 | 60394.31 | 56.8 | FTN+FLOYDS | 380 | 900 |
| 2024-03-29 | 60398.47 | 60.9 | Xinglong+BFOSC | 400 | 3000 |
| 2024-03-31 | 60400.36 | 62.8 | FTN+FLOYDS | 380 | 900 |
| 2024-04-03 | 60403.36 | 65.8 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-04-06 | 60406.36 | 68.8 | Keck+LRIS | 630/1700 | 300 |
| 2024-04-08 | 60408.28 | 70.7 | FTN+FLOYDS | 380 | 900 |
| 2024-04-11 | 60411.27 | 73.7 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-04-13 | 60413.10 | 75.6 | Bok+B&C | 700 | 900 3 |
| 2024-04-15 | 60415.50 | 78.0 | Xinglong+BFOSC | 400 | 3300 |
| 2024-04-17 | 60417.24 | 79.7 | FTN+FLOYDS | 380 | 900 |
| 2024-04-19 | 60419.26 | 81.7 | Lick+KAST | 800 | 2700 |
| 2024-04-24 | 60424.39 | 86.8 | FTN+FLOYDS | 380 | 900 |
| 2024-04-30 | 60430.24 | 92.7 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-05-02 | 60432.41 | 94.9 | FTN+FLOYDS | 380 | 1800 |
@@
Note. — Phases are reported with respect to an estimated explosion epoch of MJD 60337.55.
@@
Note. — Phases are reported with respect to an estimated explosion epoch of MJD 60337.55.

### v2

@@ 7 Summary and Conclusions
We have presented comprehensive photometric and spectroscopic data of the Type II SN 2024bch from the UV to IR over the first 150 days. The day optical spectra shows lines of H I He II, C IV, and N IV but lack C III, N III, and O IV. By day 6 all but the narrow Balmer lines persist. High-resolution spectra reveal narrow P-Cygni in H that constrains the progenitor wind to 35–40 km s-1, speeds typically associated with RSGs. Comparison between models and the early spectra indicate a mass loss rate of between . This requires a period of enhanced mass in the years to weeks before explosion as these values are four orders of magnitude higher than normal quiescent winds from RSGs.
After day 60 when the lightcurve begins to fall from the plateau we also see the emergence of multi-peaked H, likely from the SN shock still interacting with an asymmetric CSM. The lack of a corresponding red component, and the persistence of narrow H absorption points to CSM in a disc or torus that we are viewing nearly edge-on. This CSM is an order of magnitude further out than the CSM giving rise to early interaction signatures, which hints at a complex CSM geometry and mass-loss history for the progenitor of SN 2024bch. Additionally, the H I and [O I] lines all show blueshifted peaks that persist well into the nebular phase, likely from asymmetries in the SN explosion itself.
@@ Acknowledgements
We thank the anonymous referee for their constructive comments. J.E.A. is supported by the international Gemini Observatory, a program of NSF’s NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. Based on observations obtained at the international Gemini Observatory (GN-2023B-Q-135, GN-2024A-Q-403), a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation on behalf of the Gemini Observatory partnership. Time-domain research by the University of Arizona team and D.J.S. is supported by National Science Foundation (NSF) grants 2108032, 2308181, 2407566, and 2432036 and the Heising-Simons Foundation under grant #2020-1864. S.V. and the UC Davis time-domain research team acknowledge support by NSF grants AST-2008108.
@@
| 2024-01-31 | 60340.45 | 2.8 | Gemini-N+GMOS | 5800 | 600 4 |
| 2024-01-31 | 60340.46 | 2.8 | FTN+FLOYDS | 380 | 1800 |
| 2024-02-01 | 60341.48 | 3.8 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-02-03 | 60343.54 | 5.9 | FTN+FLOYDS | 380 | 900 |
| 2024-02-04 | 60344.48 | 6.8 | FTN+FLOYDS | 380 | 900 |
| 2024-02-04 | 60344.62 | 7.0 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-02-05 | 60345.46 | 7.8 | FTN+FLOYDS | 380 | 900 |
| 2024-02-05 | 60345.52 | 7.9 | LBT+PEPSI | 50000 | 5400 |
| 2024-02-06 | 60346.32 | 8.7 | FTN+FLOYDS | 380 | 600 |
| 2024-02-09 | 60349.57 | 11.9 | FTN+FLOYDS | 380 | 600 |
| 2024-02-11 | 60351.30 | 13.7 | FTN+FLOYDS | 380 | 600 |
| 2024-02-13 | 60353.30 | 15.7 | Bok+B&C | 700 | 600 3 |
| 2024-02-14 | 60354.37 | 16.7 | FTN+FLOYDS | 380 | 600 |
| 2024-02-16 | 60356.65 | 19.0 | Xinglong+BFOSC | 400 | 2100 |
| 2024-02-19 | 60359.34 | 21.7 | FTN+FLOYDS | 380 | 600 |
| 2024-02-19 | 60359.43 | 21.8 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-02-22 | 60362.40 | 24.8 | FTN+FLOYDS | 380 | 600 |
| 2024-02-25 | 60365.49 | 27.9 | FTN+FLOYDS | 380 | 600 |
| 2024-02-28 | 60368.35 | 30.7 | FTN+FLOYDS | 380 | 600 |
| 2024-02-29 | 60369.41 | 31.8 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-03-02 | 60371.10 | 33.5 | Bok+B&C | 700 | 600 3 |
| 2024-03-02 | 60371.27 | 33.6 | FTN+FLOYDS | 380 | 600 |
| 2024-03-11 | 60380.23 | 42.6 | LBT+PEPSI | 50000 | 14400 |
| 2024-03-17 | 60386.51 | 48.9 | FTN+FLOYDS | 380 | 600 |
| 2024-03-20 | 60389.30 | 51.7 | Lick+KAST | 800 | 1200 |
| 2024-03-24 | 60393.43 | 55.8 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-03-25 | 60394.31 | 56.7 | FTN+FLOYDS | 380 | 900 |
| 2024-03-29 | 60398.47 | 60.8 | Xinglong+BFOSC | 400 | 3000 |
| 2024-03-31 | 60400.36 | 62.7 | FTN+FLOYDS | 380 | 900 |
| 2024-04-03 | 60403.36 | 65.7 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-04-06 | 60406.36 | 68.7 | Keck+LRIS | 630/1700 | 300 |
| 2024-04-08 | 60408.28 | 70.6 | FTN+FLOYDS | 380 | 900 |
| 2024-04-11 | 60411.27 | 73.6 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-04-13 | 60413.10 | 75.5 | Bok+B&C | 700 | 900 3 |
| 2024-04-15 | 60415.50 | 77.9 | Xinglong+BFOSC | 400 | 3300 |
| 2024-04-17 | 60417.24 | 79.6 | FTN+FLOYDS | 380 | 900 |
| 2024-04-19 | 60419.26 | 81.6 | Lick+KAST | 800 | 2700 |
| 2024-04-24 | 60424.39 | 86.7 | FTN+FLOYDS | 380 | 900 |
| 2024-04-30 | 60430.24 | 92.6 | Gemini-N+GMOS | 930 | 300 4 |
| 2024-05-02 | 60432.41 | 94.8 | FTN+FLOYDS | 380 | 1800 |
@@
Note. — Phases are with respect to an estimated explosion epoch of MJD 60337.65.
@@
Note. — Phases are with respect to an estimated explosion epoch of MJD 60337.65.